OpenStudy (anonymous):

What is the correct solution to the equation? 8(x + 2) = 24 A. 1 B. 3 C. 5 D. 64

undeadknight26 (undeadknight26):

the solution is 1 8(x + 2) = 24 /8 /8 x + 2 = 3 - 2 -2 x = 1

OpenStudy (anonymous):

Okay. Thanks. :)